A ten sided die probability of rolling a four on 3 consecutive rolls?
The answer is 1 in 1000.
Rolling a 4 on a ten sided die is a 1 in 10 shot.
Rolling threes 4's is 1/10 x 1/10 x 1/10 which equals 1/1000.
Three coins tossed at once what the probabilty of no heads?
1/2 x 1/2 x 1/2 = 1/8 or 0.125 also 12.5%
How does one find the probability of A given B compliment?
P(A given B')=[P(A)-P(AnB)]/[1-P(B)].
In words: Probability of A given B compliment is equal to the Probability of A minus the Probability of A intersect B, divided by 1 minus the probability of B.
Throw a dice and toss a coin what is the probability that a number less than 3 and a tail result?
p(a) = 1/3, p(b) = 1/2, p(a and b) = p(a)*p(b) = 1/6
What is the probability that a person chosen at random was born on a Saturday or Sunday?
2:7
edit:
That makes the assumption that:
-birth isn't induced for various reasons - people tend to try for induced births more on weekends so that it doesn't interfere with the work week
-the ratio of people born on any given day of the week is 1:1:1:1:1:1:1
If you draw a card from a normal deck of cards what is the probabiltiy of not drawing a red jack?
Let A be the amount of cards in a normal deck that are not considered to be red. Let B be the total amount of cards in that same deck. Answer: A / B
2
10 Green marbles, 13 Blue marbles.
If two events are mutually exclusive what is the probability that both occur at the same time?
The probability is 0. Consider the event of tossing a coin . The possible events are occurrence of head and tail. they are mutually exclusive events. Hence the probability of getting both the head and tail in a single trial is 0.
What are the units of measurement for the standard error of mean?
The same units as the mean itself. If the units of the mean, are, for example miles; then the error units are miles.
Suppose a normal random variable has a mean of 72 inches and a standard deviation of 2 inches. Suppose the random variable X measures the height of adult males in a certain city. One may therefore conclude that approximately 84% of the men in this population are shorter than?
What is the probability of getting a red card from a deck of cards?
Assuming you refer to a normal deck of playing cards, either 0.481 (with jokers) or 0.5, one-half (without).
Define conditional probability with an example?
there are 3 types of conditional probability:
1. the indicate: if antecedent happens, then evidence manifests itself
example -> when tossing a coin, if it lands on tails, then you win the game
2. the subjunctive: if antecedent would happen, then evidence would manifest itself
example -> when a coin was tossed, if it would have landed on tails, i would have won the game
it is recommended for optimal bayesian inferences that your a priori distribution is indicative. If not, you could be dealing with inproper, uninformative, or hyper priors, which make decision-making and posterior determination more complex, if even possible.
Posterior distributions could very well be subjunctive. Suppose i have won the game, i could have tossed tails, but i could also have tossed heads.
What was the motivation for introducing the Gamma distribution?
It can be thought of as a generalization of the Chi-square distribution. See the link to a related WikiAnswer question below.
In tossing a coin, there are two outcomes; head or tail. Since there are two outcomes, probability of each outcome is: 0.5. Assuming a fair coin, probability of a head or tail is: 0.5. So, probability a team will win a toss is: 0.5.
Describe how the sample size affects the standard error?
Standard error (which is the standard deviation of the distribution of sample means), defined as σ/√n, n being the sample size, decreases as the sample size n increases.
And vice-versa, as the sample size gets smaller, standard error goes up. The law of large numbers applies here, the larger the sample is, the better it will reflect that particular population.
What is the probability that a couple with four children have atleast one girl?
There are 2 ways to do this problem. 1. Go to a Binomial Distribution Table where n = 4 (4 children) and P=0.5(50% probability of a girl). Probability of at least 1 girl = 1 - probability of no girls. From Binomial Distribution Table n = 0 probability is .0625. So, 1 - 0.0625 = .9375 = probability of at least 1 girl. 2. The other way is to list all the possible ways to have 4 children and count the number of ways at least 1 girl exists divided by the total number of ways to have 4 children. There are 42 ways to have 4 children, all 16 listed below: bbbb bbbg bbgb bgbb gbbb bbgg bggb ggbb gbbg gbgb bgbg bggg gggb ggbg gbgg gggg Since 15 of the 16 have at least 1 girl, the Probability of at least 1 girl = 15/16 = 0.9375, the same answer as above.
There is 2 outcomes for flipping the coin, and 6 outcomes for rolling the cube. The total outcomes for both are 2*6 = 12.
How many outcomes are there in the sample space for rolling 2 dice?
We calculate the numerator of the desired probability:
There are 6 ways any number can show up on the first die
AND
There is only 1 way that same number can show up on the second die
So since "AND" means "MULTIPLY", there are 6x1 or 6 ways
that both dice can come up the same. So the numerator of
the probability is 6.
We calculate the denominator of the desired probability:
There are 6 ways any number can show up on the first die
AND
There are 6 ways any number can show up on the second die
So since "AND" means "MULTIPLY", there are 6x6 or 36 ways
the two dice can all come up any number. So the denominator
of the probability is 36.
Probability can be defined as a ratio between the number of and number of possible outcomes?
yes it can be defined more commonly as a ratio between the number of and numbr of possible outcomes
What is the probability of getting a tail on a coin and a 6 on a dice?
probability of tails on a coin is 1/2 probablility of 6 on a die is 1/6